OpenStudy (gigi1103):

Two opposite sides of a square are increased by 5 cm to form a rectangle. If the rectangle has a perimeter of 70 cm, what was the length of a side of the original square? A. 20 cm B. 25 cm C. 15 cm D. 10 cm

OpenStudy (gigi1103):

@Warrior678 i hate to call you to help me again

OpenStudy (anonymous):

its okay

OpenStudy (gigi1103):

do you know?

OpenStudy (anonymous):

maybe give me 5min i might get the answer

OpenStudy (gigi1103):

okay

OpenStudy (anonymous):

i got 15

OpenStudy (anonymous):

ill explain it for you

OpenStudy (anonymous):

if i can

OpenStudy (gigi1103):

okay

OpenStudy (anonymous):

http://www.calculatorsoup.com/

OpenStudy (anonymous):

try this

OpenStudy (anonymous):

Two opposite sides of a square are increased by 5 cm to form a rectangle. If the rectangle has a perimeter of 70 cm, what was the length of a side of the original square? We know a square had 4 equal sides...and a rectangle has 2 equal sides and 2 equal opposite sides... if you increase 2 sides by 5...and now have a perimeter of 70...that is x + x + (x+5) + (x+5) = 70 so that is the same as 4x + 10 = 70 4x = 60 x = ?

OpenStudy (anonymous):

the answer is not 60 it is 15

OpenStudy (anonymous):

it works with most math adults or not

OpenStudy (anonymous):

for adults or not

OpenStudy (gigi1103):

mhhhhhhhhhhhhhhhhhhh think brain

OpenStudy (anonymous):

i mean it works for adults too

OpenStudy (gigi1103):

i don,t get it

OpenStudy (gigi1103):

@Here_to_Help15

OpenStudy (anonymous):

Two opposite sides of a square are increased by 5 cm to form a rectangle. If the rectangle has a perimeter of 70 cm, what was the length of a side of the original square? We know a square had 4 equal sides...and a rectangle has 2 equal sides and 2 equal opposite sides... if you increase 2 sides by 5...and now have a perimeter of 70...that is x + x + (x+5) + (x+5) = 70 so that is the same as 4x + 10 = 70 4x = 60 x = ?

OpenStudy (anonymous):

what does x = ?

OpenStudy (anonymous):

x =

OpenStudy (gigi1103):

huh? don,t we have to add five to it

OpenStudy (anonymous):

answer is 15 im 100% sure

OpenStudy (anonymous):

okay hoped i help i got to go ill fan you